# Molecular Mechanisms Underlying the Effects of Urea and the Structural Dynamics of Bovine Serum Albumin

### v1

@@ Abstract
The disruption of protein structures by denaturants like urea is well-studied, though its molecular mechanisms remain unclear. Using Molecular Dynamics (MD) simulations, we investigated how urea affects the structural stability of Bovine Serum Albumin (BSA) at concentrations from 0 M to 5 M. Our results reveal that urea induces a dehydration-rehydration cycle, characterized by displacement and partial replacement of water molecules in BSA’s hydration shell. At low concentrations, urea reduces protein-water hydrogen bonds while enhancing protein-urea interactions. At higher concentrations, urea aggregation limits these interactions, promoting rehydration and changes in tertiary structure, while secondary structure remains largely intact. These findings provide insights into the mechanisms of protein denaturation and stability by urea.
@@ I Introduction
The study of urea and other denaturing agents in aqueous protein solutions has been a central topic in biochemistry for decades. The ability of these compounds to alter the three-dimensional structure of proteins is well documented; however, the exact molecular mechanism underlying this process remains a subject of debate Wallqvist, Covell, and Thirumalai (1998); Frank and Franks (1968); Åstrand et al. (1991); Rezus and Bakker (2006); Soper, Castner Jr, and Luzar (2003); Gahtori, Gunwant, and Pandey (2023); Stumpe and Grubmüller (2007); Lindgren and Westlund (2010); Moeser and Horinek (2014); Matubayasi (2018); Nnyigide, Lee, and Hyun (2018); Kumaran and Ramamurthy (2011); Ma, Pazos, and Gai (2014); Sinibaldi et al. (2008); Monhemi et al. (2014).
@@
Several mechanisms have been proposed to explain urea-induced protein denaturation. Traditionally, two main models have been considered: the direct mechanism, where urea interacts with stabilizing intramolecular forces via hydrogen bonding and van der Waals interactions; and the indirect mechanism, where urea alters solvent properties, thereby weakening hydrophobic interactions Rossky (2008); Das and Mukhopadhyay (2009); Niether et al. (2018); Hua et al. (2008); Jha and Marqusee (2014); Guckeisen, Hosseinpour, and Peukert (2021). Recent studies, however, suggest that these mechanisms are not mutually exclusive but may act synergistically, depending on the experimental conditions Khan et al. (2019); Paladino et al. (2022); Lim, Rösgen, and Englander (2009).
The impact of urea on protein stability is influenced by multiple factors, including denaturant concentration, temperature, pH, and thermodynamic parameters Miller and Levinger (2022). Comparisons between urea and guanidinium chloride have revealed significant differences in their effects on protein conformation Baptista et al. (2023); Candotti et al. (2013); Maity et al. (2020). Computational studies indicate that urea exhibits a particular affinity for hydrophobic residues, facilitating solvation and disrupting the hydrophobic collapse Lapidus et al. (2007), ultimately leading to loss of native structural integrity Candotti et al. (2013). Recent molecular dynamics simulations have further demonstrated that, under acidic conditions, urea undergoes a localized reorientation near protonated acidic residues (Asp/Glu), acting predominantly as a hydrogen bond acceptor. However, these changes do not significantly affect urea’s interactions with the protein backbone or nonpolar residues, supporting a direct and consistent denaturation mechanism across pH conditions de Oliveira and Martínez (2020). Complementing these computational findings, surface-sensitive vibrational spectroscopy studies have experimentally confirmed that urea interacts directly with hydrophobic side chains at protein-like interfaces. This interaction reduces the hydrophobic barrier, promotes interfacial water ordering, and disrupts side chain organization, further reinforcing the direct role of urea in destabilizing the protein hydrophobic core Gahtori, Gunwant, and Pandey (2023).
@@
Additionally, molecular dynamics (MD) simulations have shown that the dipolar orientation of urea in aqueous solution can temporarily stabilize certain secondary structural elements before their complete disruption Baptista et al. (2023). This finding suggests that denaturation is not an instantaneous event but a dynamic process governed by the competition between protein-protein and protein-solvent interactions Paladino, Vitagliano, and Graziano (2023); Gooran and Kopra (2024).
@@ II Materials and methods: Systems and simulations
We investigated the impact of urea on the structural integrity of bovine serum albumin (BSA) in aqueous solutions across varying urea concentrations. The computational model utilized for urea was the Boek model Boek, Briels, and Feil (1994) with nonbonded interactions and covalent parameters sourced from the GROMOS 54A7 force field Schmid et al. (2011). For water, we employed the Simple Point Charge Extended (SPC/E) model Berendsen, Grigera, and Straatsma (1987), as developed by Berendsen and colleagues. The initial atomic coordinates for the BSA model were obtained from the crystal structure of bovine serum albumin deposited in the Protein Data Bank (PDB code 4F5S) Bujacz (2012). The topology of BSA at a standard pH of 6.5 was obtained from previous work and validated through Small-Angle X-ray Scattering (SAXS) and MD simulations Scanavachi et al. (2020). The net charge of the protein under these conditions was +11, and sodium (Na+) counterions were added to neutralize the system.
@@
The optimization and production processes of both the control system and BSA in water-urea solutions were simulated using a cubic box with periodic boundary conditions, at $300.15$ K and $1$ bar of pressure, employing a velocity rescale thermostat Bussi, Donadio, and Parrinello (2007) and Berendsen barostat Berendsen et al. (1984), respectively. The coupling time constants utilized in the thermostat and barostat were $0.1$ ps and $1.0$ ps, respectively. Electrostatic interactions were computed using the PME method Abraham and Gready (2011) with a cutoff for Van der Waals and Coulomb interactions set at $1.0$ nm. The LINCS algorithm Hess et al. (1997) was employed to constrain solute bonds. A time step of $2$ fs was maintained throughout the simulations.
@@ III.1 Hydrogen Bond Analysis
Initially, we analyzed the formation of protein-water hydrogen bonds (HBs) as a function of increasing urea concentration. For each of the eight systems corresponding to each concentration, the calculations were averaged over the final 8 ns of the simulation.
As shown in Figure 2, the presence of urea in the BSA–water system leads to a marked reduction in protein–water HBs. At 1 M urea, a decrease of approximately 25% is observed, which increases to nearly 45% as the concentration increases from 2 M to 5 M. Interestingly, beyond this point, the number of protein–water HBs plateaus at around 700, indicating that urea-induced dehydration reaches a threshold beyond which further displacement of water molecules from the protein surface is limited.
Figure 2: Total number of hydrogen bonds (HBs) formed between BSA and the components of urea:water mixtures as a function of urea concentration. Light blue bars represent the average number of HBs between the protein and water molecules, while light green bars correspond to HBs formed between the protein and urea molecules. The horizontal orange line indicates the average number of protein–water HBs in the absence of urea (0 M), used as a reference. Error bars represent the standard deviation across eight replicates. The data illustrate the progressive substitution of protein–water HBs by protein–urea HBs with increasing urea concentration, while the total number of protein–solvent HBs remains relatively constant. The supplementary material includes figures showing the HBs distributions for protein–water, protein–urea, and urea–urea interactions, normalized by the number of molecules of each species.
@@ III.2 Radial Distribution Function (RDF)
By analyzing the normalized average number of hydrogen bonds between protein and water (Figure 1 in the Supplementary Material) and its variation in the presence of urea, we show that the result is consistent with the observations in 4(a). From this, it can be observed that from 0 M to 2 M, the normalized average number of HBs between the protein and water decreases, reaching a minimum at 2 M. Beyond this minimum, the normalized average number of HBs between the protein and water increases, accompanied by the corresponding rehydration of the protein. Although the number of water molecules decreases by approximately 30% when comparing 1 M and 5 M urea concentrations (Figure 2), the normalized average number of HBs remains almost the same at both concentrations.
@@
We interpret the sustained HBs count at high urea concentrations as evidence that protein solvation by water molecules persists, while direct interactions with urea molecules diminish. Complementarily, we analyzed the formation of hydrogen bonds between the protein main chain, protein side chain, and water and urea molecules (see Figures 4 and 5 in the supplementary material).These results suggest that increasing urea concentration does not induce significant changes in the interactions between urea or water molecules and the protein main chain, but does markedly affect their interactions with the protein side chains. This supports the notion that urea–protein interactions are more likely to occur with the side chains rather than the backbone.
@@ IV Conclusion
This study employed molecular dynamics simulations to explore the effects of urea on the structural behavior of bovine serum albumin (BSA) at physiological pH. Across the range of urea concentrations tested (0–5 M), our results reveal that urea primarily induces changes at the tertiary structure level, with the secondary structure remaining largely conserved, particularly the $\alpha$-helical content. This structural integrity is consistent with previous experimental reports and highlights the stability of the N-isoform of BSA under mildly denaturing conditions Murayama and Tomida (2004); Abrosimova, Shulenina, and Paston (2016).
A key observation is the progressive dehydration of the protein surface as urea concentration increases, driven by the replacement of protein–water hydrogen bonds with protein–urea interactions. Despite this shift, the total number of hydrogen bonds involving the protein remains nearly constant, indicating a dynamic equilibrium in solvation.
Radial distribution function analysis and solvation metrics (RDF, SASA, Rg) further support a modest expansion of the protein, especially at intermediate urea concentrations, with domain-specific differences in solvent accessibility, most notably in Domain 2. This expansion correlates with subtle structural rearrangements and may reflect early stages of tertiary unfolding.
At higher concentrations, urea–urea interactions become increasingly dominant, leading to a saturation of protein–urea contacts and promoting partial rehydration of the protein. This saturation arises when all accessible interaction sites on the protein surface are occupied, preventing additional urea molecules from establishing new contacts. As a result, excess urea tends to self-associate Atahar et al. (2019), forming structured aggregates rather than interacting directly with the protein. The system thus reaches a dynamic equilibrium in which the number of protein–urea interactions no longer increases, despite the higher concentration. These findings suggest a complex, concentration-dependent interplay between dehydration, solvation, and domain flexibility, reinforcing the idea that urea-induced denaturation is a gradual and structurally selective process.
While the low number of Na+ ions in the present system limits our ability to evaluate their impact on protein destabilization, future studies could investigate how urea modulates electrostatic interactions under conditions where ionic effects are more pronounced. Previous observations in SDS micellar systems showed that urea can disrupt structure by solvating counterions and weakening electrostatic stabilization Espinosa, Grigera, and Ferrara (2018). A similar mechanism may operate in protein systems, particularly under conditions where ion–protein interactions contribute more significantly to structural integrity. Exploring these effects could provide deeper insight into the interplay between ion binding, solvation dynamics, and urea-induced denaturation.
@@ ACKNOWLEDGMENTS
We wish to thank and fondly remember Dr. Raul Grigera, who taught and shaped us in this beautiful profession, and thank the anonymous reviewers who provided thoughtful feedback on previous drafts of this article. This work was supported by grants from Consejo Nacional de Investigaciones Científicas y Tecnicas (CONICET), Agencia Nacional de Promocion Científica y Tecnologica (ANPCyT) and Universidad Nacional Arturo Jauretche (UNAJ, Argentina).

### v2

@@ Abstract
The disruption of protein structures by denaturants such as urea is well-documented, although the underlying molecular mechanisms are not yet fully understood. In this study, we employed Molecular Dynamics (MD) simulations to examine the effects of urea on the structural stability of Bovine Serum Albumin (BSA) at concentrations ranging from 0 M to 5 M. Our results reveal that urea induces a dehydration–rehydration cycle by displacing and partially substituting water molecules in BSA’s hydration shell. At lower concentrations, urea decreases protein–water hydrogen bonding while enhancing protein–urea interactions. At higher concentrations, urea tends to aggregate, which limits direct interactions with the protein, promotes rehydration, and leads to alterations in the tertiary structure, although the secondary structure remains largely preserved. These findings offer mechanistic insights into urea-induced protein denaturation and stability.
@@ I Introduction
The study of urea and other denaturing agents in aqueous protein solutions has been a central topic in biochemistry for decades. Although the ability of these compounds to disrupt protein tertiary structure is well documented, the precise molecular mechanisms underlying their action remain a subject of ongoing debate Wallqvist et al. (1998); Frank and Franks (1968); Åstrand et al. (1991); Rezus and Bakker (2006); Soper et al. (2003); Gahtori et al. (2023); Stumpe and Grubmüller (2007); Lindgren and Westlund (2010); Moeser and Horinek (2014); Matubayasi (2018); Nnyigide et al. (2018); Kumaran and Ramamurthy (2011); Ma et al. (2014); Sinibaldi et al. (2008); Monhemi et al. (2014).
@@
Multiple mechanisms have been proposed to explain urea-induced protein denaturation. Two predominant models are traditionally recognized: the direct mechanism, in which urea disrupts intramolecular stabilizing forces—such as hydrogen bonds and van der Waals interactions—by directly interacting with the protein; and the indirect mechanism, in which urea alters the structure and dynamics of the solvent, thereby weakening hydrophobic interactions that stabilize the protein’s native conformation Rossky (2008); Das and Mukhopadhyay (2009); Niether et al. (2018); Hua et al. (2008); Jha and Marqusee (2014); Guckeisen et al. (2021). Recent studies, however, suggest that these mechanisms are not mutually exclusive but may act synergistically, with their relative contributions depending on specific experimental conditions Khan et al. (2019); Paladino et al. (2022); Lim et al. (2009).
The impact of urea on protein stability depends on multiple factors, including denaturant concentration, temperature, pH, and thermodynamic parameters Miller and Levinger (2022). Comparative studies with guanidinium chloride have revealed distinct conformational effects on proteins Baptista et al. (2023); Candotti et al. (2013); Maity et al. (2020). Computational evidence indicates that urea exhibits preferential affinity for hydrophobic residues, promoting solvation and disrupting hydrophobic collapse Lapidus et al. (2007), ultimately compromising native structural integrity Candotti et al. (2013).
@@
Recent molecular dynamics simulations have shown that under acidic conditions, urea tends to reorient locally near protonated acidic residues (Asp/Glu), primarily acting as a hydrogen bond acceptor de Oliveira and Martínez (2020). Notably, these local rearrangements have minimal impact on urea’s interactions with the protein backbone or nonpolar residues, supporting the hypothesis of a largely pH-independent denaturation mechanism.Complementing these computational findings, surface-sensitive vibrational spectroscopy has experimentally confirmed urea’s direct interactions with hydrophobic side chains at protein interfaces Gahtori et al. (2023). These interactions lower the hydrophobic barrier, enhance interfacial water structuring, and disrupt side chain organization, thereby destabilizing the protein’s hydrophobic core. Additionally, molecular dynamics (MD) simulations have shown that the dipolar orientation of urea in aqueous solution can temporarily stabilize certain secondary structural elements before their complete disruption Baptista et al. (2023). This finding suggests that denaturation is not an instantaneous event but a dynamic process governed by the competition between protein-protein and protein-solvent interactions Paladino et al. (2023); Gooran and Kopra (2024).
@@ II Materials and methods: Systems and simulations
We investigated the impact of urea on the structural integrity of bovine serum albumin (BSA) in aqueous solutions across varying urea concentrations. The computational model utilized for urea was the Boek model Boek et al. (1994) with nonbonded interactions and covalent parameters sourced from the GROMOS 54A7 force field Schmid et al. (2011). For water, we employed the Simple Point Charge Extended (SPC/E) model Berendsen et al. (1987), as developed by Berendsen and colleagues. The initial atomic coordinates for the BSA model were obtained from the crystal structure of bovine serum albumin deposited in the Protein Data Bank (PDB code 4F5S) Bujacz (2012). The topology of BSA at a standard pH of 6.5 was obtained from previous work and validated through Small-Angle X-ray Scattering (SAXS) and MD simulations Scanavachi et al. (2020). The net charge of the protein under these conditions was +11, and sodium (Na+) counterions were added to neutralize the system.
@@
The optimization and production processes of both the control system and BSA in water-urea solutions were simulated using a cubic box with periodic boundary conditions, at $300.15$ K and $1$ bar of pressure, employing a velocity rescale thermostat Bussi et al. (2007) and Berendsen barostat Berendsen et al. (1984), respectively. The coupling time constants utilized in the thermostat and barostat were $0.1$ ps and $1.0$ ps, respectively. Electrostatic interactions were computed using the PME method Abraham and Gready (2011) with a cutoff for Van der Waals and Coulomb interactions set at $1.0$ nm. The LINCS algorithm Hess et al. (1997) was employed to constrain solute bonds. A time step of $2$ fs was maintained throughout the simulations.
@@ III.1 Hydrogen Bond Analysis
Initially, we analyzed the formation of protein-water hydrogen bonds (HBs) as a function of increasing urea concentration. For each of the eight systems corresponding to each concentration, the calculations were averaged over the final 8 ns of the simulation. Figure 1 in the supplementary material displays an example of the time evolution of the number of hydrogen bonds between the protein and water, averaged across the eight systems for the control condition.
As shown in Figure 2, the presence of urea in the BSA–water system leads to a marked reduction in protein–water HBs. At 1 M urea, a decrease of about 25% is observed, rising to 45% as the concentration increases from 2 M to 5 M. Interestingly, beyond this point, the number of protein–water HBs plateaus at around 700, indicating that urea-induced dehydration reaches a threshold beyond which further displacement of water molecules from the protein surface is limited.
Figure 2: Total number of hydrogen bonds (HBs) formed between BSA and the components of urea:water mixtures as a function of urea concentration. Light blue bars represent the average number of HBs between the protein and water molecules, while light green bars correspond to HBs formed between the protein and urea molecules. The horizontal orange line indicates the average number of protein–water HBs in the absence of urea (0 M), used as a reference. Error bars represent the standard deviation across eight replicates. The data illustrate the progressive substitution of protein–water HBs by protein–urea HBs with increasing urea concentration, while the total number of protein–solvent HBs remains relatively constant. Figures 2, 3, and 4 in supplementary material show the HBs distributions for protein–water, protein–urea, and urea–urea interactions, normalized by the number of molecules of each species.
@@ III.2 Radial Distribution Function (RDF)
By analyzing the normalized average number of hydrogen bonds between protein and water (Figure 2 in the Supplementary Material) and its variation in the presence of urea, we show that the result is consistent with the observations in 3(a). From this, it can be observed that from 0 M to 2 M, the normalized average number of HBs between the protein and water decreases, reaching a minimum at 2 M. Beyond this minimum, the normalized average number of HBs between the protein and water increases, accompanied by the corresponding rehydration of the protein. Although the number of water molecules decreases by approximately 30% when comparing the systems at 1 M and 5 M urea concentrations (Figure 2), the normalized average number of hydrogen bonds (HBs) remains nearly constant at both concentrations.
@@
We interpret the sustained HBs count at high urea concentrations as evidence that protein solvation by water molecules persists, while direct interactions with urea molecules diminish. Complementarily, we analyzed the formation of hydrogen bonds between the protein main chain, protein side chain, and water and urea molecules (see Figures 5 and 6 in the supplementary material).These results suggest that increasing urea concentration does not induce significant changes in the interactions between urea or water molecules and the protein main chain, but does markedly affect their interactions with the protein side chains. This supports the notion that urea–protein interactions are more likely to occur with the side chains rather than the backbone.
@@ IV Conclusion
This study employed molecular dynamics (MD) simulations to investigate the effects of urea on the structural behavior of bovine serum albumin (BSA) at physiological pH. Across the tested urea concentrations (0-5 M), our results demonstrate that urea primarily disrupts the tertiary structure, while the secondary structure-particularly the $\alpha$-helical content-remains largely preserved. This structural resilience is consistent with previous experimental findings and highlights the stability of BSA’s N-isoform under mildly denaturing conditions Murayama and Tomida (2004); Abrosimova et al. (2016).
A key observation is the progressive dehydration of the protein surface with increasing urea concentration, driven by the replacement of protein–water hydrogen bonds with protein–urea interactions. Despite this shift, the total number of hydrogen bonds involving the protein remains nearly constant, suggesting a dynamic solvation equilibrium.
Analysis of radial distribution functions (RDF) and solvation metrics-including solvent-accessible surface area (SASA) and radius of gyration (Rg)-reveals modest protein expansion, particularly at intermediate urea concentrations. Domain-specific differences in solvent accessibility were observed, with the most pronounced changes occurring in Domain 2. This expansion correlates with subtle structural rearrangements, which may represent early stages of tertiary structure unfolding.
At higher concentrations, urea-urea interactions become dominant, saturating protein-urea contacts and promoting partial rehydration of the protein. This saturation occurs when all accessible surface sites on the protein are occupied, preventing further urea binding. As a result, excess urea molecules self-associate to form structured aggregates rather than interacting with the protein Atahar et al. (2019). The system thus reaches a dynamic equilibrium, wherein protein-urea interactions plateau despite further increases in urea concentration. These findings highlight a concentration-dependent balance between dehydration, solvation, and domain-specific flexibility, supporting the notion that urea-induced denaturation progresses gradually and selectively affects distinct aspects of protein structure.
Although the limited number of Na+ ions in our system constrains the evaluation of their role in protein destabilization, future studies should investigate how urea modulates electrostatic interactions under more physiologically relevant ionic conditions. Previous studies on SDS micellar systems have shown that urea can disrupt structural organization by solvating counterions and weakening electrostatic stabilization Espinosa et al. (2018). A similar mechanism may be operative in protein systems, particularly in cases where ion–protein interactions play a significant role in maintaining structural integrity. Exploring these effects could yield novel insights into the interplay between ion binding, solvation dynamics, and urea-induced protein denaturation.
@@ ACKNOWLEDGMENTS
We wish to thank and fondly remember Dr. Raúl Grigera, who taught and shaped us in this beautiful profession. We also thank the anonymous reviewers for their thoughtful feedback on earlier drafts of this article. This work was supported by grants from the Consejo Nacional de Investigaciones Científicas y Técnicas (KE3-11220210100918CO) and the National University Arturo Jauretche (80020230100029UJ).
The Supplementary Material accompanying this study includes detailed methodological descriptions, auxiliary computations, and extended analyses that support the findings and conclusions presented in the main manuscript. Specifically, it contains an analysis of the equilibrium conditions of the studied systems and an evaluation of the number of hydrogen bonds formed between the protein and water or urea molecules, along with their respective normalizations. In addition, it reports the number of hydrogen bonds between urea molecules and their normalization. Finally, it presents an analysis of the average formation of hydrogen bonds involving the protein main chain, protein side chain, and water molecules. The material is provided in a structured and accessible format to facilitate reference.